# Spacetime games subsume causal contextuality scenarios

### v3

@@ 5 Concrete impact of the mapping on the simplicity of constructs and proofs
We now turn to a few examples showing that looking at a causal contextuality scenario via its associated spacetime game is directly useful and simplifies certain constructs, reasonings, and proofs.
@@ 5.1 Good covers
In the paper introducing causal contextuality scenarios [8], it is conjectured that certain covers dubbed “good covers” characterize certain causal contextuality scenarios for which desirable properties apply—such as the sheaf property. We first show that these good covers are straightforward to define from the mapping provided in this paper; in the next section, we will show that it is also straightforward to prove that the sheaf property applies in the case of pure strategies.
Concretely, causal contextuality scenarios that have a good cover can be defined as the range of the mapping of Definition 7. Causal contextuality scenarios with a cover that is not a good cover can still be mapped to an alternating spacetime game, however, this mapping does not roundtrip and the correspondence has no semantic or physical meaning in this case.
Definition 9.
A causal contextuality scenario is said to have a good cover if it is in the range of the mapping defined in Definition 7.
Theorem 5.
The mapping of Definition 7 becomes bijective if we consider its corestriction to causal contextuality scenarios with good cover.
Proof.
Replacing the codomain of an injective function with the range of the function makes it bijective. ∎
It is also useful to know that one can change the cover of any causal contextuality scenario to be a good cover, keeping its other components intact.
Theorem 6.
Given a causal contextuality scenario, it is always possible to find a causal contextuality scenario with a good cover that only differs from the first one in the cover.
Proof.
Given a causal contextuality scenario, one can:
  • Replace its cover with the powerset of $X$.
  • Map it to an alternating spacetime game.
  • Map back to a causal contextuality scenario, which will have a good cover by construction.
Note that the powerset of $X$ is not necessarily a good cover, so that the final causal contextuality scenario’s good cover need not be the powerset of $X$. The suchly obtained causal contextuality scenario only differs from the initial causal contextuality scenario in the cover, as the mapping leaves the measurements, outcomes, and enabling relations intact. This is because (i) measurements correspond to Alfred’s information sets, left-hand sides of enabling relations to Bob’s nodes, and outcomes to Alfred’s actions at a given information set, (ii) since we start with the full powerset of $X$, no measurement, outcome or enabling relation is lost while mapping the causal contextuality scenario to a spacetime game (as could occur more generally with some covers that are not good covers), (iii) by definition, the mapping the spacetime game to the final causal contextuality scenario with good cover also does not lose any measurement, outcome or enabling relation. ∎
Note that the good cover is not generally unique given measurements, outcomes, and enabling relations. The determination of the most sensible cover to model a specific experiment requires additional knowledge about the considered Hilbert spaces, observables, and commutation properties.
5.2 Global strategies
@@
If we view a causal contextuality scenario as its corresponding spacetime game, this game has a corresponding strategic form as we explained in 2020 [10]: it is the (reduced) strategic form of the associated game in extensive form with imperfect information, which was constructed 75 years ago by Harold Kuhn [12][9]. A pure strategy in this strategic form provides the desired global strategy function.
5.3 The sheaf of strategies
Let us take any causal contextuality scenario with good cover.
In the flat case, it was shown by Abramsky and Brandenburger [21] that the presheaf of strategies has the sheaf property because it also involves the gluing of partial functions on discrete space that agree on their overlaps. Thanks to the spacetime game framework, it is straightforward to translate this argument to any causal contextuality scenario with good cover.
Indeed, if we take a pure strategy as defined by Kuhn instead of the form given by Abramsky et al [8] in terms of game semantics, we note that these strategies are partial functions on a discrete space ($\mathcal{I}$) to actions. Thus, we can generally consider the restrictions of these partial functions to smaller domains, from which it follows as in shown by Abramsky and Brandenburger [21] that the suchly defined presheaf of strategies has the sheaf property.
This was left as an open question by [8] to determine “good covers” for which this is the case, and we have thus shown that Definition 9 successfully addresses this open question. The simplicity of the construction of the strategy sheaf of Nature demonstrates the expressive power of the spacetime game framework, itself building on 75 years of game theory literature.
5.4 Mixed strategies and empirical models
@@
In the special case of flat games (no adaptive measurements), it was shown by Abramsky and Brandenburger [21] that the existence of a deterministic hidden variable model realizing an empirical model is implied by the existence of a global section for it. With spacetime games, this is intuitive to understand: we observe that a mixed strategy of nature is the same mathematical object as a global section of the event distribution presheaf and we remember that spacetime games have a natural cover given by the projection of the support of their complete histories to nature.
@@
Given a causal contextuality scenario with a good cover and an empirical model, the existence of a global section that is consistent with an empirical model implies the existence of a deterministic hidden variable model for this empirical model.
@@
We have just generalized the result of Abramsky and Brandenburger [21] from the flat case to any causal contextuality scenario with a good cover. The reason for the short proof is not that it is trivial; rather, it builds on previous work including Kuhn’s construct. This is consistent with our qualitative argument given by Baczyk and Fourny [6] that Nashian game theory is not compatible with quantum physics. The formulation in terms of presheaves and the sheaf property [21] strengthens this argument.
@@ 5.5 Perfect information
If we do not require perfect information for spacetime games, this introduces another physical justification for imperfect information in the resulting extensive form, namely, noncontextuality because non-singleton information sets in the spacetime game (which lead to the fusion of the corresponding information sets in the extensive form) arise from the different contexts in which a measurement can be carried out, with the result independent of the context. This is also consistent with the fact that locality is a special case of non-contextuality.
@@ 5.6 Temporal contextuality scenarios
Temporal contextuality scenarios were introduced by Searle et al [20] as a special case of causal contextuality scenario in which the measurements are timelike-separated. If we consider the associated spacetime games, temporal contextuality scenarios correspond to games in extensive form with perfect information. Furthermore, the lookbacks—corresponding to limited memory on past experiments— can be modeled in the spacetime game framework with imperfect information, i.e., by connecting identical measurements carried out in distinct pasts together. All the above simplified constructs apply: causal secured covers, global strategies, the presheaf of pure strategies, and the link between the existence of a global section of an empirical model and the existence of a deterministic hidden variable model.
@@ 6 Related work
Hardy models [14], PR boxes [15], and GHZ models [16] are representable with spacetime games.
The framework also maps to contextuality by default [23], which corresponds to having perfect information. Forcing random variables corresponding to measurement outcomes to have the same value formally corresponds to grouping the corresponding nodes into the same information set.
Sheaves of sections as constructed by Gogioso and Pinzani [13] correspond to a special case of spacetime games with perfect information in which measurements are performed independently of previous measurement outcomes, without merging the information sets (see also Section 6.b. in [8]), and with unique causal bridges. The fact that the sheaf property obtains is consistent with our results, which generalize the case covered by Gogioso and Pinzani [13] to all alternating spacetime games, and causal contextuality scenarios that have a natural cover.
A definition of free choice is given by Renner and Colbeck [24], as a decision (called spacetime variable) being independent of anything it could not have caused. A deterministic resolution algorithm for all spacetime games with perfect or imperfect information, and which circumvents impossibility theorems (such as the sheaf property of the strategy distribution presheaf not generally holding) by weakening free choice, is given in [5], with the special case of games in normal form [25] and of games in extensive form with perfect information [17]. A high-level discussion of the non-Nashian paradigm in the context of quantum physics interpreted as a game between the observer and nature is also available [26].
The seeding philosophy of the non-Nashian approach as well as the notion of causal bridge is attributable to Dupuy [18] (counterfactual decision theory) and the game theoretic seed and initial conjectures were made by [19].
@@
[31] give an excellent argument of why impossibility theorems do not disprove the existence of local theories. Hance and Hossenfelder [32] also are of the opinion that free choice can be weakened. Wharton and Argaman [33] refer to this possibility as (non-)future-input-dependent parameters, and it is also known as measurement independence (see for example the work of Storz et al [34] for the explicit mention of this assumption in Bell experiments) or setting independence [35].
7 Conclusion
We would like to conclude by sharing our scientific belief that only spacetime games with perfect information correspond to a physical reality, and that spacetime games with imperfect information are, above all, a useful mathematical tool for reductio ad absurdum proofs that nature is contextual. This is aligned with work by Dzhafarov [23] and on keeping locality as a fundamental principle of nature, provided we weaken measurement independence and assume non-Nashian free choice instead [6].
@@ 8 Acknowledgements
I thank my students at ETH who contributed to the game framework in various projects over the years: Felipe Sulser, Emilien Pilloud, Ramon Gomm, Luc Stoffer, Martin Kucera, Michael Baczyk, Florian Pauschitz.

### v4

@@ 4 Discussion
We now turn to a few examples showing that looking at a scenario via its associated spacetime game is directly useful and simplifies certain constructs, reasonings, and proofs.
@@ 4.1 Good covers
In the paper introducing causal contextuality scenarios [11], it is conjectured that certain covers dubbed “good covers” (or “causally secured covers” in subsequent talks) characterize certain causal contextuality scenarios for which desirable properties apply—such as the sheaf property. We are not aware, at the time of writing, of a subsequent publication giving a formal definition of these good covers, but show further down that if we define causally secured covers as in this paper, they provide such a candidate in which the sheaf property holds for the (pure) strategy presheaf [^17] But not to empirical models, which correspond to mixed strategies..
4.2 The pure strategy sheaf of a spacetime game
@@
If we view a causal contextuality scenario as its corresponding spacetime game, this game has a corresponding strategic form as we explained in 2020 [10]: it is the strategic form of the associated game in extensive form with imperfect information, which was constructed 75 years ago by Harold Kuhn [13][14]. A pure strategy in this strategic form provides the desired global strategy function.
Worded differently, Kuhn’s strategic form can be used to obtain the presheaf of (pure) strategies similar to a presheaf of events whose global sections are the pure strategies of the (non-reduced) strategic form. The reduced strategic form eliminate some redundancies by grouping pure strategies in equivalence classes, however, the reduced pure strategies would then no longer be global sections.
In the flat case, it was shown by Abramsky and Brandenburger [12] that the presheaf of events (which is the flat counterpart to the presheaf of strategies) has the sheaf property because it involves the gluing of partial functions on discrete space that agree on their overlaps. Thanks to the spacetime game framework, it is straightforward to translate this argument to any acyclic causal contextuality scenarios with unique causal bridges and a causally-secured cover.
If we define the pure strategy presheaf as indicated above, then we note that these strategies are partial functions on a discrete space ($\mathcal{I}_{A}=X$) to actions. Thus, we can generally consider the restrictions of these partial functions to smaller domains, from which it follows as in shown by Abramsky and Brandenburger [12] that the suchly defined presheaf of strategies has the sheaf property.
This was left as an open question by [11] to determine “good covers” for which this is the case, and we have thus shown that Definition 11 successfully addresses this open question at least for acyclic scenarios with unique causal bridges. The simplicity of the construction of the strategy sheaf of Nature demonstrates the expressive power of the spacetime game framework, itself building on 75 years of game theory literature.
Since the categories Game and Scenario are equivalent, one can also instead define the strategy presheaf of a spacetime game by applying the definition given in [11] to the equivalent scenario. We conjecture that this construction is structurally equivalent to the pure strategy presheaf defined above.
4.3 The mixed strategy presheaf of a spacetime game
An original contribution of [11] is the generalization of empirical models to scenarios involving causality. Empirical models are obtained through the composition of a distributions functor with the (pure) strategy sheaf. This can be done on different kinds of semirings leading to (signed or unsigned) probabilistic or possibilistic semantics.
@@ 4.3 The mixed strategy presheaf of a spacetime game
Thus, a corollary insight of this paper is that applying the distribution functor to the (pure) strategy sheaf [11] is akin to introducing mixed strategies as distributions over pure strategies as is classically done in game theory. The (pre)sheaf structure then additionally involves restrictions of pure or mixed strategies to smaller supports or, the other way around, asks whether a compatible set of distributions can be glued back into a global section. The existence of a global section corresponds to the case in which a mixed strategy of Nature successfully explains the obtained correlations with various choices of contexts.
Thus, a natural synonym for the distribution presheaf obtained by composing the distribution functor with the sheaf of pure strategies would be presheaf of mixed strategies.
@@
In the special case of flat games (no adaptive measurements), it was shown by Abramsky and Brandenburger [12] that the existence of a deterministic hidden variable model realizing an empirical model is implied by the existence of a global section for it. With spacetime games, this is intuitive to understand: we observe that a mixed strategy of nature is the same mathematical object as a global section of the event distribution presheaf and we remember that spacetime games have a natural cover given by the facets of the projection of the support of their complete histories to nature.
@@
Given an alternating spacetime game and an empirical model, the existence of a mixed strategy (global section) that is consistent with an empirical model implies the existence of a deterministic hidden variable model for this empirical model.
@@
We have just generalized the result of Abramsky and Brandenburger [12] from the flat case to acyclic causal contextuality scenarios with a causally-secured cover and unique causal bridges. The reason for the short proof is not that it is trivial; rather, it builds on previous work including Kuhn’s construct. This is consistent with our qualitative argument given by Baczyk and Fourny [8] that Nashian game theory is not compatible with quantum physics. The formulation in terms of presheaves and the sheaf property [12] strengthens this argument.
@@ 4.4 Perfect information
If we do not require perfect information for spacetime games, this introduces another physical justification for imperfect information in the resulting extensive form, namely, noncontextuality because non-singleton information sets in the spacetime game (which lead to the fusion of the corresponding information sets in the extensive form) arise from the different contexts in which a measurement can be carried out, with the result independent of the context. This is also consistent with the fact that non-locality is a special case of contextuality.
@@ 4.5 Temporal contextuality scenarios
Temporal contextuality scenarios were introduced by Searle et al. [24] as a special case of causal contextuality scenario in which the measurements are timelike-separated. If we consider the associated spacetime games, temporal contextuality scenarios correspond to games in extensive form with perfect information. Furthermore, the lookbacks—corresponding to limited memory on past experiments— can be modeled in the spacetime game framework with imperfect information, i.e., by connecting identical measurements carried out in distinct pasts together. All the above simplified constructs apply: causally-secured covers, pure strategies, the (pure) strategy sheaf, and the link between the existence of a global section of an empirical model (a mixed strategy) and the existence of a deterministic hidden variable model.
4.6 Information matrices
Thanks to the equivalence of categories, and since causal contextuality scenarios adapt and repurpose concrete domains and information matrices by Kahn and Plotkin [26], we gain the interesting insight that spacetime games are directly related to this earlier contribution, where cells correspond to nodes of a spacetime game, that is, $\Gamma$ corresponds to $\mathcal{N}$, the set of values $V$ corresponds to the set of actions $\mathcal{A}$, $\mathcal{V}$ corresponds to $\chi$, and the enabling function $\mathcal{E}$ corresponds to the labeled incoming edges of a node ($\mathcal{R}$ and $\sigma$). Further even earlier related work to information matrices includes that of Nielsen, Plotkin and Winskel [27]. Thus, the equivalence of categories connects our work to a dense body of literature, which could lead to further future insights combining the two research directions.
@@ 5 Related work
Hardy models [17], PR boxes [18], and GHZ models [19] are representable with spacetime games. The framework also maps to contextuality by default [28], which corresponds to having perfect information. Forcing random variables corresponding to measurement outcomes to have the same value formally corresponds to grouping the corresponding nodes into the same information set.
Sheaves of sections as constructed by Gogioso and Pinzani [16] correspond to a special case of spacetime games with perfect information in which measurements are performed independently of previous measurement outcomes, without merging the information sets (see also Section 6.b. in [11]), and with unique causal bridges. The fact that the sheaf property obtains is consistent with our results, which generalize the case covered by Gogioso and Pinzani [16] to all alternating spacetime games, and the causal contextuality scenarios with no cycles, unique causal bridges, and a causally-secured cover.
A definition of free choice is given by Renner and Colbeck [5] and Sengupta [3], as a decision (called spacetime variable) being independent of anything it could not have caused. A deterministic resolution algorithm for all spacetime games with perfect or imperfect information, and which circumvents impossibility theorems (such as the sheaf property of the strategy distribution presheaf not generally holding) by weakening free choice, is given in [7], with the special case of games in normal form [29] and of games in extensive form with perfect information [20]. A high-level discussion of the non-Nashian paradigm in the context of quantum physics interpreted as a game between the observer and nature is also available [30].
The seeding philosophy of the non-Nashian approach as well as the notion of causal bridge is attributable to Dupuy [21] (counterfactual decision theory) and the game theoretic seed and initial conjectures were made by [22]. The work of Petri [31] on causally secured events around the same time is independently related to Dupuy’s work, and Petri nets, dating back to as early as 1939 in their first form, can also be seen as an early precursor to the concept of causal bridge in a special-relativistic sense.
@@
[36] give an excellent argument of why impossibility theorems do not disprove the existence of local theories. Hance and Hossenfelder [37] also are of the opinion that free choice can be weakened. Wharton and Argaman [38] refer to this possibility as (non-)future-input-dependent parameters, and it is also known as measurement independence (see for example the work of Storz et al. [39] for the explicit mention of this assumption in Bell experiments) or setting independence [40].
6 Conclusion and future work
We would like to conclude by sharing our scientific belief that only acyclic causal contextuality scenarios with unique causal bridges and causally-secured covers correspond to a physical reality, and that more general scenarios (with cycles, or non-unique causal bridges, or more general covers) are, above all, a useful mathematical tool for reductio ad absurdum proofs that nature is contextual.
Whether scenarios with cycles can exist is the subject of a broad debate in the quantum foundations community known as indefinite causal order. For example, see an argument against the existence of such scenarios in [41].
Non-unique causal bridges physically mean that a measurement carried out under specific enabling conditions is physically the same measurement as a measurement carried out under different enabling conditions. What this means is closely related to the difference between the classical, Nashian view of free choice as opposed to the non-Nashian approach. Indeed, the core difference between the two approaches is to what extent the past light cone of a decision deterministically determines this decision: the link is more loose in the Nashian view, while it is tight in the non-Nashian view. In the non-Nashian view, which is the direction of our research programme, the tight link between a decision (by an observer or the universe) and the past light cone is consistent with unique causal bridges: a measurement carried out in specific enabling conditions is a not the same as a measurement (counterfactually) carried out in different conditions.
Finally, causally-secured cover pertain to locality as can be seen from the criteria used in our definition: saying that the support of the set of events on the left-hand side of an enabling condition must be included in every context in which the right-hand side appears is consistent with the semantics of enabling conditions. The maximality of a causally-secured cover can be interpreted in terms of a form of non-signaling, in the sense that choices made at some location cannot influence the available contexts at a remote location, and the local cover restrictions are also a feature of locality.
This discussion is aligned with work by Dzhafarov [28] and on keeping locality as a fundamental principle of nature, provided we weaken measurement independence and assume non-Nashian free choice instead [8].
Future work involves a generalization of the categorical equivalence between games and causal contextuality scenarios to a more limited set of assumptions. It also involves further exploration on the consequences of the
@@ 7 Acknowledgements
I thank my students at ETH who contributed to the game framework in various projects over the years: Felipe Sulser, Emilien Pilloud, Ramon Gomm, Luc Stoffer, Martin Kucera, Michael Baczyk, Florian Pauschitz. I thank Ben Moseley, Renato Renner, Kuntal Sengupta, Ognyan Oreshkov, Jacopo Surace, Beatrix Hiesmayr, Marek Szopa, Kim Vallée, Stefan Weigert for stimulating discussions and my co-author Michael Baczyk for his valuable feedback on the mapping between games and scenarios.